# Ravenea: A Benchmark for Multimodal Retrieval-Augmented Visual Culture Understanding

### v1

@@ 7 Conclusion
We introduce Ravenea, a novel benchmark dataset designed to comprehensively evaluate the cultural sensitivity of diverse multimodal retrievers across 14 SOTA VLMs. The benchmark consists of culturally contextualized queries and image captions from eight countries, curated and paired with Wikipedia passages ranked by human annotators based on their cultural relevance. Our findings highlight the potential of RAG for visual cultural understanding, particularly when lightweight VLMs are enhanced with culturally-aware multimodal retrievers such as CaCLIP, which consistently outperform their non-augmented counterparts – with the caveat that the largest models in each VLM family often exhibit diminishing returns when integrated with RAG. Notably, the inclusion of culture-sensitive questions during data annotation significantly improves effectiveness of multimodal retrievers and enhances performance in downstream tasks.
8 Acknowledgment
We would like to thank all the annotators for their work. Special thanks to Nico Lang, Peter Ebert Christensen, Zhaochong An, Stella Frank, Srishti Yadav, for providing helpful research advice. Serge Belongie and Jiaang Li are supported by the Pioneer Centre for AI, DNRF grant number P1. Ivan Vulić is supported by a personal Royal Society University Research Fellowship ‘Inclusive and Sustainable Language Technology for a Truly Multilingual World’ (no 221137). Wenyan Li is supported by the Lundbeck Foundation (BrainDrugs grant: R279-2018-1145).
Appendix A Limitations
While Ravenea establishes a solid foundation for advancing the study of visual culture understanding with retrieval augmentation, it has three limitations that warrant future attention. First, due to budgetary constraints, the dataset’s cultural scope is currently limited to eight countries and eleven categories. Although this selection introduces meaningful diversity, it does not comprehensively represent the global spectrum of cultural perspectives, particularly those of underrepresented or marginalized communities. Second, our use of Wikipedia as the primary external knowledge source introduces inherent biases and may lack the depth, plurality, and contextual richness necessary for nuanced cultural interpretation. Finally, due to resource limitations, we were unable to include certain proprietary VLMs that require paid APIs, such as Gemini 2.5 Pro [52] and Claude Opus 3.7 [53]. We hypothesize that their performance would be comparable to GPT-4.1, which was included in our evaluation, but this remains an open empirical question.
Appendix B Future Directions
Our work opens several avenues for advancing visual culture understanding in multimodal models. First, expanding Ravenea to include more countries, cultural categories, and diverse knowledge sources, beyond Wikipedia, would improve coverage and reduce institutional bias. Second, future benchmarks could include richer tasks beyond cVQA and cIC, such as culture-grounded object recognition, historical retrieval, and symbolic interpretation, to better capture cultural semantics. Third, our results suggest a need for culturally-aware evaluation metrics, particularly for text generation. The limited effectiveness of retrieval augmentation in larger models also warrants further study, especially regarding how cultural knowledge is integrated and utilized. Together, these directions aim to support the development of more culturally-sensitive and globally robust vision-language models.
Appendix C Ethics Statement
@@
This work focuses on improving the cultural awareness of VLMs through retrieval-augmented methods. All data used in the construction of the Ravenea benchmark were sourced from publicly available datasets and Wikipedia, a community-curated open-access knowledge base. To protect individual privacy, we apply automated face detection and blur all identifiable faces in images prior to release. To mitigate cultural bias and ensure broad representation, the benchmark includes images and documents spanning eight countries and eleven cultural domains, curated and annotated by a diverse group of annotators. We provide detailed documentation of the annotation process and guidelines to support transparency and reproducibility. While enhancing cultural understanding is a central goal of this work, we acknowledge that culture is inherently complex, dynamic, and context-dependent. Consequently, the benchmark cannot capture the full richness of any cultural context. Finally, this work does not involve any personally identifiable information (PII), biometric data, or sensitive attributes. All human annotators were compensated fairly, and data annotation adhered to ethical guidelines for responsible research.
Appendix D Experimental Details for Multi-modal Retrievers
Dataset.
We integrate responses from three annotation questions per data point into a continuous scale ranging from $-3$ to $3$, where higher values indicate stronger cultural relevance. To ensure fair evaluation across regions, we adopt the train-validation-test split strategy from class-imbalance loss [54], yielding an approximate 85-5-10 data split. This setup guarantees that validation and test sets contain a balanced set of images per country (Figure 8), thereby mitigating evaluation bias from skewed geographic distributions. During training, we employ random cropping for images as a sample-level augmentation to normalize the distribution of training instances across countries, further mitigating region-specific sampling biases.
@@ Appendix D Experimental Details for Multi-modal Retrievers
Hyperparameters VisualBERT VL-T5 LLaVA-OneVision-7B CLIP-L/14@224px SigLIP2-SO/14@384px
@@
lr warmup ratio - 0.1 - - -
weight decay - - - - -
Max Epoch 100 20 10 50 50
Patience 10 10 3 5 5
early stopping Yes
@@
In this work, we adopt different sets of hyperparameters as VisualBERT, VL-T5, LLaVA-OneVision-7B, CLIP-L/14@224px, SigLIP2-SO/14@384px. For VisualBERT, VL-T5, and LLaVA-OneVision-7B we follow the setting in [44, 45, 46]. We show the training hyperparameters in rerankering experiments for all models in Table 5. All experiments are conducted using a maximum of 2 Nvidia H100 GPUs.
Appendix E Data Statistics
Table 6: Statistics of images in each country.
India Mexico Indonesia China Nigeria Korea Spain Russia
408 349 309 303 223 151 142 77
Table 7: Statistics of images in each category.
Architecture Cuisine History Art Daily Life Companies Sports & Recreation Transportation Religion Nature Tools
403 402 278 275 185 80 73 68 52 31 21
We then apply cryptographic hashing (SHA-256) to identify and remove duplicate images, resulting in a cleaner and more distinct cultural image set. Consequently, Ravenea comprises images collected from eight countries across four continents, spanning eleven distinct categories. The distribution of images by country and by category is detailed in Tables 6 and 7, respectively.
Appendix F Evaluation of VLMs
For closed-weight model, GPT-4.1, we directly call the corresponding API. For open-source models, we use vllm [55]. During the evaluation, to ensure the stability of the results, we set the temperature parameter to 0.0 and the maximum output length to 256. All open-weight models are listed in Table 8. To avoid the impact of the length the retrieved content, we use the first 256 words in the top-1 Wikipedia document.
@@ Appendix F Evaluation of VLMs
LLaVA-OneVision-7B [19] llava-hf/llava-onevision-qwen2-7b-ov-hf
Phi-4-Multimodal [21] microsoft/Phi-4-multimodal-instruct
Pixtral [20] mistral-community/pixtral-12b
Qwen2.5VL family [23]
@@
DeepSeek-VL2 family [25]
@@
OpenGVLab/InternVL3-38B
Gemma3 family [22]
@@
Appendix G Annotation Details
Based on the initial BM25 retrieval results, we refine the cultural relevance label of retrieved documents via human annotation. We found that directly asking annotators to rate overall cultural relevance on a continuous scale (e.g., 0–10) led to unreliable and inconsistent labels. This difficulty arises from several factors: (1) the semantic meaning of intermediate scores is ambiguous, (2) annotators tended to overemphasize a few salient visual elements [56], and (3) small numerical differences (e.g., between 5 and 6) often fail to reflect meaningful distinctions, especially given the cognitive load of processing lengthy Wikipedia documents, resulting in intra-annotator variance even on repeated examples. Instead, given an image–caption–document triplet, we decomposed cultural relevance into three interpretable and independently verifiable dimensions: country association, topic alignment, and explicit visual representation.
Prior to the annotation process, all annotators are required to carefully review a detailed instruction file outlining the relevance criteria and annotation guidelines. To ensure proper understanding of the guidelines, annotators are required to complete a mock annotation test and correctly answer all questions before proceeding with the actual annotation tasks. For each image-caption pair, annotators are presented with the top 10 Wikipedia documents retrieved by BM25. They are asked to assess whether each article provides meaningful background or contextual information that is directly relevant to the cultural elements described in the caption or depicted in the image.
Appendix H A New Metric: RegionScore
To evaluate the extent to which captions reference specific geopolitical regions, we introduce a RegionScore. This metric measures whether a caption contains explicit references to a country or its common demonyms. $N$ be the total number of samples. $\text{PRED}_{i}$ denote predicted captions for the $i$-th sample. $C_{i}$ be the country name associated with the $i$-th sample. $\text{Adj}(C_{i})$ denote the set of adjectives or demonyms associated with $C_{i}$. $\mathcal{T}_{i}=\{C_{i}\}\cup\text{Adj}(C_{i})$ denote the set of region-related terms for the $i$-th sample. We define binary indicators for each sample:
$\delta^{\text{pred}}_{i}=\begin{cases}1&\text{if any }t\in\mathcal{T}_{i}\text{ appears in }\text{PRED}_{i}\\ 0&\text{otherwise}\end{cases}$ (5)
The RegionScore for predicted captions are then computed as:
$\text{RegionScore}_{\text{PRED}}=\frac{1}{N}\sum_{i=1}^{N}\delta^{\text{pred}}_{i}$ (6)
These scores reflect the proportion of captions that include explicit regional identifiers. A higher score indicates stronger region-awareness in the captioning. For the ground truth in CCUB dataset, the $\text{RegionScore}_{\text{GT}}=99\%$.
Appendix I Details of Human Evaluation in the cIC Task
We randomly sample 10 images and generate captions using 14 vision-language models (VLMs) under three configurations: CaCLIP, CLIP, and a no-retrieval baseline, yielding 420 captions in total. Four expert annotators participated in the evaluation. For each image, they were presented with caption triplets—one per retrieval configuration—and asked to assess approximately 35 such triplets each. Annotators are instructed to select the caption that most accurately and appropriately reflects the cultural context depicted. To evaluate annotation consistency, we randomly sample 30 triplets and assign them to a fifth expert annotator. Inter-annotator agreement (IAA) between this annotator and the original four, measured using Cohen’s $\kappa$, is 0.595.
Appendix J Additional Results
@@ Appendix I Details of Human Evaluation in the cIC Task
Figure 9: Performance of the 14 VLMs equipped with CaCLIP in cVQA task.
As illustrated in Figures 9 and 10, the models exhibit diverse cultural preferences. Notably, most models achieve relatively stronger performance on Chinese and Indian cultural contexts in the cVQA task, and on Chinese and Korean contexts in the cIC task.
@@
Figure 10: Performance of the 14 VLMs equipped with CaCLIP in cIC task.
We report the CIDEr, ROUGE-L, BERTScore, and CLIPScore metrics for 14 VLMs on the cIC task in Table 9.

### v2

@@ 6 Conclusion
We presented Ravenea, a new benchmark for multimodal retrieval-augmented visual culture understanding that enables the evaluation of how external cultural knowledge improves VLMs. By pairing culture-centric visual question answering and culture-informed captioning with 11,396 human-ranked Wikipedia documents across eight countries and eleven categories, Ravenea provides the first large-scale testbed for studying multimodal retrieval in culturally grounded settings. Our experiments across seven multimodal retrievers and seventeen VLMs show that incorporating cultural grounding consistently improves both retrieval quality and downstream task performance. Notably, lightweight VLMs benefit significantly from culture-aware retrieval, highlighting multimodal retrieval augmentation as an efficient way to enhance culturally situated understanding and reasoning. Overall, Ravenea offers a standardized framework for diagnosing limitations and advancing research in multimodal retrieval, grounding, and generation for VLMs in culture domain. We hope this benchmark will drive future work toward more culture-sensitive and globally inclusive multimodal intelligence.
Ethics Statement
@@
This work focuses on improving the cultural awareness of VLMs through retrieval-augmented methods. All data used in the construction of the Ravenea benchmark were sourced from publicly available datasets and Wikipedia, a community-curated open-access knowledge base. To protect individual privacy, we apply automated face detection and blur all identifiable faces in images prior to release. To mitigate cultural bias and ensure broad representation, the benchmark includes images and documents spanning eight countries and eleven cultural domains, curated and annotated by a diverse group of annotators. We provide detailed documentation of the annotation process and guidelines to support transparency and reproducibility. While enhancing cultural understanding is a central goal of this work, we acknowledge that culture is inherently complex, dynamic, and context-dependent. Consequently, the benchmark cannot capture the full richness of any cultural context.
Unavoidably, when models fail, errors may go beyond reduced accuracy, leading to misattributed traditions or distorted cultural practices. In heritage contexts, such failures may undermine trust in digital tools or inadvertently disrespect the communities represented. Reliance on broad but uneven sources like Wikipedia can further compound these risks by amplifying existing biases and privileging cultures with richer digital footprints, echoing broader challenges in AI for cultural heritage (Münster et al., 2024). Such distortions could affect heritage management by marginalizing underrepresented voices, eroding local agency, or encouraging overreliance on automated outputs. To mitigate these risks, we emphasize safeguards such as human-in-the-loop validation with cultural experts, transparent documentation of dataset limitations, and fairness-oriented evaluation protocols. Looking forward, retrieval grounded in domain-specific knowledge bases curated by GLAM (Galleries, Libraries, Archives, and Museums) institutions offers a path toward higher-quality cultural data. We also highlight the importance of systematically incorporating expert human evaluation to validate both cultural relevance and interpretive accuracy.
Finally, this work does not involve any personally identifiable information (PII), biometric data, or sensitive attributes. All human annotators were compensated fairly, and data annotation adhered to ethical guidelines for responsible research.
Acknowledgment
We would like to thank all the annotators for their work. Special thanks to Nico Lang, Peter Ebert Christensen, Zhaochong An, Stella Frank, Srishti Yadav, for providing helpful research advice. This project is mainly supported by the National Research Foundation Singapore under the AI Singapore Programme (AISG Award No: AISG3-RPGV-2025-016). Jiaang Li and Serge Belongie are supported by the Pioneer Centre for AI, DNRF grant number P1. Ivan Vulić is supported by a personal Royal Society University Research Fellowship ‘Inclusive and Sustainable Language Technology for a Truly Multilingual World’ (no 221137). Wenyan Li is supported by the Lundbeck Foundation (BrainDrugs grant: R279-2018-1145). Yang Deng is support by the Lee Kong Chian Fellowship awarded by Singapore Management University.
Reproducibility Statement
Section 2 and Appendix E, F and L details the data collection and processing steps. Experimental setups are provided in Appendix I. Code and dataset are available here.
Appendix A The usage of Large Language Models (LLMs)
We use LLMs as writing aids, limited to polishing grammar, improving clarity, and refining readability. Moreover, during data curation, we used LLMs to generate culture-informed captions solely as auxiliary references for annotators; these LLM outputs were not directly included in the final dataset. All technical content, results, and conclusions are original and authored entirely by the listed contributors. The responsibility for the manuscript’s scientific validity and integrity rests fully with the authors.
Appendix B Limitations
While Ravenea establishes a solid foundation for advancing the study of visual culture understanding with retrieval augmentation, it has three limitations that warrant future attention. First, due to budgetary constraints, the dataset’s cultural scope is currently limited to eight countries and eleven categories. Although this selection introduces meaningful diversity, it does not comprehensively represent the global spectrum of cultural perspectives, particularly those of underrepresented or marginalized communities. Second, our use of Wikipedia as the primary external knowledge source introduces inherent biases and may lack the depth, plurality, and contextual richness necessary for nuanced cultural interpretation. Finally, due to resource limitations, we were unable to include certain proprietary VLMs that require paid APIs, such as Gemini 2.5 Pro (Team et al., 2023) and Claude Opus 3.7 (Anthropic, 2024). We hypothesize that their performance would be comparable to GPT-4.1 (Achiam et al., 2023), which was included in our evaluation, but this remains an open empirical question.
Appendix C Future Directions
Our work opens several avenues for advancing visual culture understanding in multimodal models. In constructing Ravenea, we algorithmically generated a raw dataset spanning 30 countries, comprising over 3,900 images and 78,000 documents (Table 4). Due to budget constraints, annotations were limited to data from eight countries. Therefore, extending Ravenea to include more countries, cultural categories, and diverse knowledge sources, beyond Wikipedia, would improve coverage and reduce institutional bias. Second, future benchmarks could include richer tasks beyond cVQA and cIC, such as culture-grounded object recognition, historical retrieval, and symbolic interpretation, to better capture cultural semantics. Third, our results suggest a need for culturally-aware evaluation metrics, particularly for text generation. The limited effectiveness of retrieval augmentation in larger models also warrants further study, especially regarding how cultural knowledge is integrated and utilized. Exploring retrieval pipelines grounded in domain-specific knowledge bases curated by GLAM (Galleries, Libraries, Archives, and Museums) institutions could provide higher-quality cultural data and more reliable grounding. Finally, systematic human evaluation by cultural experts remains indispensable for validating both relevance and interpretive accuracy. Together, these directions aim to support the development of more culturally-sensitive and globally robust vision-language models.
Appendix D Data Statistics
Table 4: Statistics of the Ravenea dataset. The dataset is constructed by curating existing sources and augmenting them with wiki-derived documents to broaden cultural knowledge coverage and enhance content diversity.
Dataset Countries Images Documents Pairs Questions Captions
CVQAselected 7 1,213 8,194 12,130 2,331 -
CCUBselected 5 655 4,354 6,550 - 655
Ravenea 8 1,868 11,396 18,680 2,331 655
BM25 filtered 30 3,912 78,240 - - -
Table 5: Statistics of images in each country in Ravenea.
India Mexico Indonesia Nigeria China Korea Spain Russia
408 347 309 223 214 148 142 77
Table 6: Statistics of images in each category in Ravenea.
Architecture Cuisine Daily Life History Art Companies Sports & Recreation Transportation Religion Nature Tools
403 402 347 278 86 80 73 68 52 31 21
We then apply cryptographic hashing (SHA-256) to identify and remove duplicate images, resulting in a cleaner and more distinct cultural image set. Consequently, Ravenea comprises images collected from eight countries across four continents, spanning eleven distinct categories. The distribution of images by country and by category is detailed in Tables 5 and 6, respectively. During constructing Ravenea, we have constructed a comprehensive, algorithmically generated raw dataset covering 30 countries, comprising over 3,900 images and 78,000 documents (see Table 4). Owing to budgetary limitations, for Ravenea, we have only annotated data from eight of these countries. Overall, we believe that both the raw algorithmically generated data and our human-labeled Ravenea could be a valuable source for improving the cultural understanding in VLM tasks.
Appendix E Annotation Details
Based on the initial BM25 retrieval results, we refine the cultural relevance label of retrieved documents via human annotation. We found that directly asking annotators to rate overall cultural relevance on a continuous scale (e.g., 0–10) led to unreliable and inconsistent labels. This difficulty arises from several factors: (1) the semantic meaning of intermediate scores is ambiguous, (2) annotators tended to overemphasize a few salient visual elements (Chen et al., 2015), and (3) small numerical differences (e.g., between 5 and 6) often fail to reflect meaningful distinctions, especially given the cognitive load of processing lengthy Wikipedia documents, resulting in intra-annotator variance even on repeated examples. Instead, given an image–caption–document triplet, we decomposed cultural relevance into three interpretable and independently verifiable dimensions: country association, topic alignment, and explicit visual representation.
Figure 8: Statistics on the Yes / No response distributions for the annotation questions.
Prior to the annotation process, all annotators are required to carefully review a detailed instruction file outlining the relevance criteria and annotation guidelines. To ensure proper understanding of the guidelines, annotators are required to complete a mock annotation test and correctly answer all questions before proceeding with the actual annotation tasks. For each image-caption pair, annotators are presented with the top 10 Wikipedia documents retrieved by BM25. They are asked to assess whether each article provides meaningful background or contextual information that is directly relevant to the cultural elements described in the caption or depicted in the image. The statistics of the annotations are presented in Figure 8.
Table 7: Statistics of missing documents, along with the positive and negative ratios for each country.
Country Positive & Negative Ratio(%) Missing Ratio(%)
India 54.42 / 45.58 9.56
Russia 54.42 / 45.58 0.0
Spain 48.17 / 51.83 0.0
Nigeria 34.47 / 65.53 0.45
Mexico 38.37 / 61.63 25.55
Indonesia 41.26 / 58.74 0.32
Korea 69.59 / 30.41 19.59
China 56.12 / 43.88 10.75
As shown in Table 7, regions such as Mexico (25.55%) and Korea (19.59%) exhibit notably high missing ratios, underscoring the indispensable role of human expertise in identifying culturally specific documents that purely automated retrieval methods fail to capture. This missing ratio serves as an informative diagnostic metric: higher values directly reflect the effectiveness of our human-in-the-loop pipeline in enriching the dataset with highly local and otherwise overlooked content, ultimately attesting to the superior cultural coverage and quality of Ravenea. Furthermore, Table 7 reports the distribution of positive versus negative labels for document relevance across regions, offering additional insight into cultural representation variance.
Appendix F Quality Control
To evaluate the quality of the GPT-4o-generated captions, we randomly sampled 50 captions from the subset of Chinese images, and employ two local checkers to check the quality in Table 8. Each annotator assessed the accuracy of the captions, and we report their individual accuracies along with the IAA to measure consistency. Beyond objective correctness, we also asked annotators to provide subjective ratings (5-point Likert scale) of caption quality along two dimensions: naturalness (fluency and readability) and completeness (coverage of visual content). The results further show the GPT-4o caption is suitable for the next step to retrieve the Wikipedia documents.
Table 8: Generated caption quality check. IAA denotes inter-annotator agreement.
Acc. (Annotator 1) Acc. (Annotator 2) IAA Naturalness (1–5) Completeness (1–5)
92% 94% 0.85 4.98 4.66
To ensure the quality and consistency of our annotations, we implement several quality control methods. First of all, prior to the annotation process, all annotators are required to carefully review a detailed instruction file outlining the relevance criteria and annotation guidelines. To ensure proper understanding of the guidelines, annotators are required to complete a mock annotation test and correctly answer all questions before proceeding with the actual annotation tasks (see Figure 13). We also perform an additional quality check on a subset of the dataset. Specifically, for each selected countries, we employ an additional local quality checker who is tasked with manually reviewing the annotations to verify their accuracy and adherence to the guidelines. The quality checker reviews a random sample of annotated items, focusing on both the relevance labels and the justification behind any edge cases, such as borderline relevance or use of the “Cannot be determined” label. If inconsistencies or deviations from the annotation guidelines are identified, the affected samples are flagged for re-annotation. The overall acceptance rate from the meta quality checkers is 98.2%. The inter-annotator agreement (IAA) Cohen’s Kappa ($\kappa$) between the meta checker and annotator on the sampled annotations is 0.83.
Appendix G Correlation between Automatic Metrics and Human Judgments
To assess the correlation between human preferences and automatic evaluation metrics, we compute Kendall’s $\tau$ rank correlation. Human annotations are segmented according to the output chunks produced by corresponding 14 VLMs (no Qwen3VL family) with each retriever variant (Ravenea-CLIP-based, CLIP-based, and non-RAG). Within each segment, we calculate the selection winning ratio for each retrieval method, yielding a human preference vector formed by concatenating these ratios across all evaluation instances. For the automatic evaluation, we extract BERTScore, CIDEr, ROUGE-L, and CLIPScore for each corresponding retrieval variant and VLM. These scores are similarly concatenated into a metric-based vector. Finally, we compute Kendall’s $\tau$ between the human and metric vectors to quantify the consistency between automatic rankings and human judgments.
Appendix H Details of Human Evaluation in the cIC Task
We randomly sample 10 images and generate captions using 14 VLMs under three configurations: Ravenea-CLIP, CLIP, and a no-retrieval baseline, yielding 420 captions in total. [^7] The variants of the Qwen3-VL models we used were not released at the time of this evaluation. Four expert annotators participated in the evaluation. For each image, they were presented with caption triplets—one per retrieval configuration—and asked to assess approximately 35 such triplets each. Annotators are instructed to select the caption that most accurately and appropriately reflects the cultural context depicted. To evaluate annotation consistency, we randomly sample 30 triplets and assign them to a fifth expert annotator. Inter-annotator agreement (IAA) between this annotator and the original four, measured using Cohen’s $\kappa$, is 0.595.
Appendix I Implementation Details
I.1 Dataset.
@@ Appendix I Implementation Details
We integrate responses from three annotation questions per data point into a continuous scale ranging from 3-3 to $3$, where higher values indicate stronger cultural relevance. To ensure fair evaluation across regions, we adopt the train-validation-test split strategy from class-imbalance loss (Cui et al., 2019), yielding an approximate 85-5-10 data split. We also performed strict de-duplication on the training and validation sets, guaranteeing that no identical documents appear across the training, validation, testing splits. This setup guarantees that validation and test sets contain a balanced set of images per country in per task (Figure 9), thereby mitigating evaluation bias from skewed geographic distributions. During training, we employ random cropping for images as a sample-level augmentation to normalize the distribution of training instances across countries, further mitigating region-specific sampling biases.
I.2 Multimodal Retrievers.
We finetune CLIP-L/14@224px (Radford et al., 2021), SigLIP2-SO/14@384px (Tschannen et al., 2025). We also finetune a VisualBERT-based (Li et al., 2019; 2020) reranker following standard BERT-style setups (Nogueira and Cho, 2019), and adapt two multimodal generative models, such as VL-T5 (Cho et al., 2021) and LLaVA-OneVision-7B (Li et al., 2025), for end-to-end document retrieval (Chen et al., 2022; Yuan et al., 2024; Sun et al., 2023). We integrate responses from three annotation questions per data point into a continuous scale ranging from 3-3 to $3$, where higher values indicate stronger cultural relevance.
@@
Hyperparameters VisualBERT VL-T5 LLaVA-OV-7B CLIP-L/14@224px SigLIP2-SO/14@384px
@@
lr warmup ratio - 0.1 - 0.03 0.03
Max Epoch 100 20 10 2 2
Patience 10 10 3 - -
early stopping Yes Yes Yes No No
@@
In this work, we adopt different sets of hyperparameters as VisualBERT, VL-T5, LLaVA-OV-7B, CLIP-L/14@224px, SigLIP2-SO/14@384px. For VisualBERT, VL-T5, and LLaVA-OV-7B we follow the setting in (Chen et al., 2022; Yuan et al., 2024; Sun et al., 2023). We show the training hyperparameters in rerankering experiments for all models in Table 9. All experiments are conducted using a maximum of 2 Nvidia H100 GPUs.
I.3 Vision-language models
@@
LLaVA-OneVision-7B (Li et al., 2025) llava-hf/llava-onevision-qwen2-7b-ov-hf
Phi-4-Multimodal (Abouelenin et al., 2025) microsoft/Phi-4-multimodal-instruct
Pixtral (Agrawal et al., 2024) mistral-community/pixtral-12b
Qwen2.5VL family (Bai et al., 2025b)
@@
Qwen3VL family (Bai et al., 2025a)
Qwen/Qwen3-VL-2B-Instruct
Qwen/Qwen3-VL-8B-Instruct
Qwen/Qwen3-VL-32B-Instruct-FP8
DeepSeek-VL2 family (Wu et al., 2024)
@@
OpenGVLab/InternVL3-78B-AWQ
Gemma3 family (Team et al., 2025)
@@
We benchmark open and closed-weight widely used VLMs on Ravenea, leveraging various retrievers against non-RAG baselines, assessing retrieval effectiveness across models of different sizes. The open-sourced models include LLaVA-OneVision-7B (Li et al., 2025), Pixtral-12B (Agrawal et al., 2024), Phi-4 Multimodal-Instruct (Abouelenin et al., 2025), Gemma3-4B-Instruct and 27B-Instruct (Team et al., 2025), Qwen2.5-VL-Instruct (3B, 7B, 72B) (Bai et al., 2025b), Qwen3-VL-Instruct(3B, 7B, 32B), InternVL3 (2B, 8B, 78B) (Zhu et al., 2025), and Deepseek-VL2 variants (Tiny and Base) (Wu et al., 2024). For the closed models, we adopt GPT-4.1 (Achiam et al., 2023) (accessed on 2025/04/14). [^8] Knowledge cutoff: June 1, 2024; https://platform.openai.com/docs/models/gpt-4.1. For proprietary model, GPT-4.1, we directly call the corresponding API. For open-sourced models, we use vllm (Kwon et al., 2023). During the evaluation, to ensure the stability of the results, we set the temperature parameter to 0.0 and the maximum output length to 2048. All open-weight models are listed in Table 10. To avoid the impact of the length the retrieved content, we use the first 256 words in the top-1 Wikipedia document.
Appendix J More Results
Figure 10: Average improvement across 16 open-source VLMs for different countries with 2 retrievers on both cIC RegionScore and cVQA accuracy.
@@
(a) Performance of the 16 VLMs equipped with Ravenea-CLIP in cVQA task.
(b) Performance of the 16 VLMs equipped with Ravenea-CLIP in cIC task.
@@
Figure 11: Performance of the 16 VLMs equipped with Ravenea-CLIP across tasks and countries.
As illustrated in Figures 11(a) and 11(b), the models exhibit diverse cultural preferences. Notably, most models achieve relatively stronger performance on Indian cultural contexts in both cIC and cVQA tasks. As shown in Figures 10, VLMs demonstrate varying degrees of performance shifts across cultural contexts in both the cIC and cVQA tasks. On cIC, Ravenea-CLIP yields noticeable improvements over their original counterparts in Korea, India and Nigeria, indicating the effectiveness of retrieval adaptation. On cVQA, while Ravenea-CLIP generally exhibit large performance improvement compared to non-RAG baselines, certain countries, such as Indonesia, experience exhibit diminishing returns.
Table 11: Ablation at cultural context. Models in gray are frozen retrievers, noisy characters, and unrelated documents. VLMs augmented with finetuned retriever generally perform better.
Retriever Average DeepSeek-VL2-Tiny Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-78B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OV-7B $cVQA_{Accuracy\uparrow}$ W/O RAG 66.0 49.8 62.7 52.6 83.3 64.1 71.3 84.2 66.5 79.4 42.6 73.2 62.7 Unrelated documents 60.6 46.4 58.9 53.6 76.1 60.8 67.0 82.3 59.8 74.6 41.1 63.2 43.5 Random Characters 62.3 44.0 57.9 52.6 79.4 63.2 69.9 83.7 54.1 75.1 41.1 68.9 57.4 CLIP-L/14@224px 69.4 50.2 63.6 66.0 82.8 73.7 78.5 84.2 69.4 76.6 67.0 76.6 44.5 Ravenea-CLIP (ours) 71.2 50.7 64.1 69.9 82.3 76.1 81.3 85.2 69.9 79.9 69.4 75.1 50.2 $cIC_{RegionScore\uparrow}$ W/O RAG 54.4 26.5 53.1 69.4 71.4 38.8 49.0 65.3 73.5 75.5 28.6 53.1 49.0 Unrelated documents 48.8 18.4 46.9 57.1 61.2 28.6 51.0 63.3 59.2 71.4 24.5 51.0 53.1 Random Characters 51.2 24.5 44.9 61.2 67.3 36.7 44.9 67.3 61.2 73.5 28.6 59.2 44.9 CLIP-L/14@224px 64.1 38.8 59.2 71.4 69.4 65.3 65.3 71.4 73.5 73.5 61.2 69.4 51.0 Ravenea-CLIP (ours) 67.7 46.9 61.2 77.6 75.5 69.4 67.3 73.5 75.5 75.5 63.3 71.4 55.1